OpenStudy (anonymous):

3k + 2j^2 k^2 = 5 make j the subject

OpenStudy (asadkarim7):

5 - 3k = 2j^2 k^2 take J^2 common J^2 ( 2k^2 ) = 5 -3k j ^2 = 5 - 3k / 2k^2 j = sqrt 5-3k / 2k^2